OpenStudy (anonymous):

PLEASEEEEE Find the vertex of the parabola formed by the following quadratic function: y = -x2 + 10x - 1 (-10, -201) (5, 24) (-5, -76) (10, -1)

OpenStudy (btaylor):

the formula for the x-coordinate of the vertex of a parabola (ax2 + bx + c) is -b/2a. so here, b=10 and a = -1 10/-2 = -5, so C is the answer.

OpenStudy (kainui):

An easier way, is to take the derivative and set it equal to 0 and solve for the x coordinate. y=x^2+10x-1 y'=2x+10 0=2x+10 x=-5 plug this into the original equation to find the y-coordinate.